1. Lerato is earning R18 per hour making pancakes.how much is she earning in cents per second?
2. Moosa can wash 6 Windows in 21 minutes. How many minutes will it takes him to wash 10 Windows?
1
Expert's answer
2020-10-08T15:44:01-0400
  1. One cent is a a hundredth of R , so R 1 = 100 c. Every hour Lerato earns "R\, 18 = 18 \cdot 100 \, c= 1800 \,c". Now we need to find how much cents she is earning per second. One hour is 60 minutes, every minute is 60 seconds. So there is 3600 s in 1h.

Lerato earns

"\displaystyle 1800\frac{c}{h} = 1800 \cdot \frac{c}{3600 s} = \frac{1800}{3600} \frac{c}{s} = \frac{1}{2}\frac{c}{s}"

Answer: 1/2 or 0.5 cent per second.

  1. Moosa can wash 6 Windows in 21 minutes. How much time does he spend for one window? This is "\displaystyle \frac{21}{6}" minutes per window. To wash 10 windows, he needs

"\displaystyle 10 \cdot \frac{21}{6} = \frac{21 \cdot 10}{6} = \frac{21 \cdot 5}{3} = 7 \cdot 5 =35" minutes.

Answer: 35 minutes.